# Galaxy Morphologies at Cosmic Noon with JWST: A Foundation for Exploring Gas Transport with Bars and Spiral Arms

### v1

@@ 4.4 Interacting/Mergers
The fraction of galaxies classified as interacting or merging systems shows a clear decline with cosmic time, decreasing from 14.2% in the higher redshift bin ($2.52.0<z<$ 2.5) to 5.9% in the lower redshift bin ($1.30.8<z<$ 1.3). This trend aligns with the expected decrease in merger activity at lower redshifts, driven by declining galaxy densities and interaction rates (e.g., Conselice, 2007; Lin et al., 2008; Lotz et al., 2008; Rodriguez-Gomez et al., 2015). Despite our modest sample size, this decline remains consistent with theoretical and observational predictions in the framework of the hierarchical growth model in which major mergers become less frequent at later cosmic times.
@@ 5 Quantitative morphological metrics
We performed a quantitative analysis of the morphological properties of the sample to complement the visual classification. To achieve this, we employed several established non-parametric morphological indicators commonly utilized in morphological studies to characterize the structural features of the galaxies in our sample. These metrics include the Concentration, Asymmetry, and Smoothness (CAS) statistics (Conselice, 2003), the Multinode, Intensity, and Deviation (MID) statistics (Freeman et al., 2013), and the Gini and $M_{20}$ coefficients (Lotz et al., 2004). All measurements are quantified using statmorph (Rodriguez-Gomez et al., 2019). Here is a brief description of them.
@@
MID: The Multinode, Intensity and Deviation (MID) statistics allow for the identification of disturbed morphologies, such as mergers or interacting systems. Multinode ($M$) measures the presence of multiple peaks in the light distribution, indicating complex or multi-component structures. Intensity ($I$) assesses the brightness contrast between these peaks and their surroundings, highlighting areas of intense star formation or disruptions. Deviation ($D$) quantifies the departure of the light distribution from a smooth, symmetric model, further aiding in the detection of structural disturbances.
@@
To assess the robustness of the morphological indicators, we conducted a systematic visual inspection of diagnostic images produced by starmorph (e.g., Figure 4 in Rodriguez-Gomez et al., 2019). The reliability of each measurement was assessed based on the following criteria: How well the model replicates the galaxy’s light distribution (e.g., the agreement between the observed and model images); quality of the residual image, where well-fit galaxies exhibit minimally structured residuals and poor fits show significant leftover flux or systematic artefacts; convergence of fitted parameters, ensuring they fall within physically reasonable ranges; and absence of contamination from neighbouring sources or image artefacts that could bias measurements. The evaluation of these diagnostic images, along with assessments of segmentation maps and PSF matching, were used to ensure the accuracy of the non-parametric morphological measurements. After applying this filtering process, we identified a high-quality subset of 1,222 galaxies that met our reliability criteria. We adopt this subset for the results and analysis presented in this work.
Figure 13: Notable parameter spaces where the combination of morphological metrics separate well different morphologies. In each panel, the red line corresponds to the line of best separation, which corresponds to the minimal contamination $\mathcal{C}_{\mathrm{min}}$ (defined in Appendix C), and the light red area indicates other possible line positions which achieve contamination within 0.01 of the minimal contamination, i.e., $C\in(\mathcal{C}_{\mathrm{min}},\mathcal{C}_{\mathrm{min}}+0.01)$. Left: Gini vs $M_{20}$ plane for galaxies with and without bulges. The line $G=-0.52M_{20}-0.4$, separates the main cluster of galaxies with bulges (71%) and those without bulges (74%). Middle: Gini vs $M_{20}$ plane for galaxies identified as interacting/mergers. The line $G=-0.29M_{20}+0.07$ (red) shows a systematic separation of the cluster of datapoints in this parameter space, with 50% of visually classified mergers and 93% of non-mergers correctly separated. For comparison, we show with the black dashed line the boundary line established by Lotz et al. (2008) and discuss the difference in the main text. Right: Gini vs Asymmetry $A$ plane for clumpy and non-clumpy galaxies. The separation line is given by $G=0.229A+0.448$, with 81.8% of clumpy galaxies below it and 64.1% of the non-clumpy galaxies above.
@@ 5.1 Dependence on redshift and stellar mass
Our findings also agree with trends reported in the literature, particularly in Yao et al. (2023) and Ren et al. (2024), which similarly noted stable morphological indicators across redshifts but a clear dependence on galaxy stellar mass. Figure 12 illustrates these comparisons, showing the mean values reported by Yao et al. (2023) for a F [^4] W-matched sample in the redshift range $30.8<z<$ 3, as well as the median values from Ren et al. (2024) for $30.5<z<$ 3. To enable a consistent comparison, we reverse-applied the observational corrections used in Ren et al. (2024) to their data, which accounted for effects such as noise and finite PSF resolution, aligning their values with our F [^4] W-matched dataset. In both cases, we find a strong agreement between the trends in our analysis and those reported by these studies.
@@ 5.2 Connection with visual classification
With the extensive set of morphological parameters computed in this analysis, we systematically explore how different galaxy classifications separate across various parameter spaces derived from statmorph . Specifically, we identify parameter combinations that effectively distinguish galaxies with and without bulges, mergers and non-mergers, and clumpy versus non-clumpy galaxies. To quantitatively assess the separation between classifications, we compute a separation efficiency metric based on misclassification rates. The methodology for defining and calculating this metric, including the optimization of classification boundaries, is detailed in Appendix C, where we provide a full set of separation efficiencies for each classification. Given the wealth of possible parameter combinations, we highlight only the most informative and physically meaningful ones, prioritizing those that offer the clearest distinctions between morphological classes that have also been studied in the literature. However, all possible combinations, including those not well studied in other studies are presented in Tables 4-6.
We emphasize that, with JWST’s unprecedented resolution, wavelength coverage, and depth, widely used morphological metrics, originally designed for lower-resolution datasets such as HST imaging, must be carefully reassessed. However, in this study, we focus on identifying broad, top-level trends in our sample rather than proposing precise recalibrations of these classification boundaries. A robust recalibration, which is beyond the scope of this work, should be carried out using larger statistical samples that fully leverage the growing wealth of high-quality JWST data, as well as the extensive galaxy catalogues from ongoing large surveys produced with EUCLID, which will provide millions of galaxies at local and intermediate redshifts (e.g., Euclid Collaboration et al. 2025a, b).
Bulges: For galaxies visually identified as containing bulges, the Gini–$M_{20}$ plane provides one of the strongest separations between bulge-dominated and non-bulge galaxies (left panel of Figure 13). The optimal separation line, given by $G=-0.52M_{20}-0.41$, correctly separates 74% of galaxies without bulges, as shown by the clear systematic offset between the two populations in the figure. This offset aligns with previous studies (e.g., Conselice, 2003; Lotz et al., 2008), which demonstrated that non-parametric morphology indicators are effective in distinguishing bulge-dominated galaxies. The correlation between bulge presence, high Gini values, and low $M_{20}$ values supports the interpretation that bulge-dominated galaxies have centrally concentrated light distributions and less fragmented substructures, making them easier to identify in this parameter space. Specifically, low $M_{20}$ values indicate that a galaxy’s brightest pixels are concentrated near the centre, which is characteristic of systems with prominent bulges.
@@
Mergers: Several parameter combinations effectively distinguish interacting and merging galaxies, including the Gini–$M_{20}$ plane, MID statistics, and CAS parameters. The full separation efficiencies for mergers and non-mergers are presented in Table 5 in the Appendix.
@@
In the Gini–$M_{20}$ plane, we define a separation line that maintains a merger fraction of $\approx 10\%$, matching both our measured fraction and the findings of Lotz et al. (2008). The resulting threshold, given by $G=-0.29M_{20}+0.07$, correctly captures 93% of non-mergers (middle panel of Figure 13). For comparison, we show the boundary line estimated by a dedicated study of mergers and morphology Lotz et al. (2008), which analyzed $\approx 3,000$ galaxies with HST imaging at $1.20.2<z<$ 1.2 and found a separation boundary of $G=-0.14M_{20}+0.33$. We find notable differences between our classification and theirs, particularly a 0.15 slope difference in the separation line. Moreover, our separation achieves a contamination value of 0.57, while adopting the Lotz et al. boundary in our dataset would result in a higher contamination value of 0.74. The differences in our boundary lines likely stem from differences in sample selection, imaging depth, rest-frame wavelength, spatial resolution and classification objectives. Whereas Lotz et al. optimized their boundary specifically for merger detection, we focused on a wider range of morphologies. Despite these differences, the overall separation trends remain similar, reinforcing that Gini vs $M_{20}$ remains a useful discriminator and is effective, at least in a global sense, within our sample.
@@
Among the CAS statistics, the asymmetry parameter provides another useful merger discriminator. Figure 14 (top panel) illustrates the separation, where visually identified mergers are skewed toward higher asymmetry values, while non-mergers peak at lower asymmetries. The separation threshold that maintains a merger fraction of 10% closely aligns with the empirical merger criterion proposed by Conselice (2003) (dashed black), reinforcing the utility of asymmetry as a merger diagnostic.
@@
Similarly, the MID statistics also effectively separate mergers from non-mergers. As shown in the bottom panels of Figure 14, the Multinode (M), Intensity (I), and Deviation (D) distributions exhibit systematic differences between the two populations. Interacting systems consistently show higher values across all three statistics, reinforcing the idea that MID metrics capture morphological disturbances and asymmetries associated with mergers and their irregular light distributions.
@@
Clumps: The Gini coefficient is particularly effective in distinguishing clumpy from non-clumpy galaxies, especially when combined with other morphological metrics. Among these, the Gini–$A$ (asymmetry) plane provides the clearest separation (right panel of Figure 13), highlighting the strong correlation between light concentration and structural irregularity in clumpy systems. The complete set of separation efficiencies for clumpy galaxies is provided in Table 6 in the Appendix.
Figure 14: Top panel: Distribution of the Asymmetry statistics ($A$ in $CAS$) in our sample, which shows a clear systematic separation around value $A=0.23$. Maintaining a merger fraction of $\approx 10\%$, we obtain the vertical red line $A=0.332$ that separates 44% of mergers and 93% of non-mergers, with only a minor offset from the boundary line of $A=0.35$ introduced by Conselice (2003). Bottom three panels: The Multinode, Intensity and Deviation (MID) statistics reveal systematic differences in distributions between visually classified mergers and non-mergers, as expected, given their design to capture morphological features of disturbed systems.
@@ 6 Morpho-Kinematic Correlations
For this discussion, we focus on global morpho-kinematic trends on a population-wide angle rather than detailed substructures, examining how the identified morphologies relate with dominant dynamical support as measured by $v/\sigma$. We use the homogenized $v_{\mathrm{rot}}(r_{\mathrm{eff}})$ and $\sigma_{0}$ values obtained from various subsamples outlined in §2.3. Our analysis reveals a notable distinction in the dynamical properties of galaxies classified as disks versus non-disks. For the 155 galaxies with kinematics measurements, 125 ($\approx 81\%$) were visually classified as disks, while 30 ($\approx 19\%$) were not identified as disks.
@@
The histogram on the right of Figure 15 presents the $v/\sigma$ distributions for visually classified disks and non-disks, where we examined their $v/\sigma$ ratios relative to thresholds of 1, 3 and 10, commonly used in literature, corresponding to varying degrees of rotational support (e.g. Genzel et al., 2006; Kassin et al., 2012; Wisnioski et al., 2015; Wisnioski et al., 2019; Stott et al., 2016; Tiley et al., 2019; Gillman et al., 2020; Förster Schreiber & Wuyts, 2020). A high $v/\sigma$ ratio suggests a dynamically cold, rotation-dominated system typical of disk galaxies, whereas a low $v/\sigma$ ratio indicates a dynamically hot, dispersion-dominated system signalling, e.g., more spheroidal geometries or stronger gravitational perturbations in merging systems.
Figure 15: Scatter plots of $v_{\mathrm{rot}}(r_{\mathrm{eff}})/\sigma_{0}$ as a function of effective radius (left) and Sérsic index (right) for disk (orange circles) and non-disk (blue squares) galaxies. The histograms above each panel show the normalized distributions of effective radius and Sérsic index, while the histogram on the right represents the distribution of $v_{\mathrm{rot}}(r_{\mathrm{eff}})/\sigma_{0}$. The red horizontal lines indicate reference values at $v_{\mathrm{rot}}(r_{\mathrm{eff}})/\sigma_{0}=1,3,$ and $10$.
@@
Among the 125 visually classified disk galaxies, 88 (70%) have $v/\sigma>3$, indicating strong rotational support. The remaining 37 disk galaxies (30%) have $v/\sigma<3$, suggesting either weaker rotational support or the presence of non-circular motions (pure turbulence but also radial motions contributing to higher $\sigma_{0}$ in modelling of the typically seeing-limited IFU data included here). For non-disk galaxies, the trend is reversed. Only 12 galaxies (40%) have $v/\sigma>3$, while the majority, 18 galaxies (60%), have $v/\sigma<3$. This distribution reflects the expectation that non-disk galaxies exhibit lower rotational support, with their dynamics often dominated by random motions rather than ordered rotation. In the case of marginal rotational support of $v/\sigma=1$, only 3 galaxies identified as disks and 5 from the other categories are below this boundary.
@@
The apparent correlation between disk morphology and high $v/\sigma$ is reassuring, yet it is important to consider observational limitations in the kinematics measurements that could affect the correlation. One such effect is beam smearing, which artificially enhances velocity dispersions in smaller galaxies (e.g., Newman et al., 2013). In particular, compact galaxies at high redshift may exhibit inflated velocity dispersions due to unresolved velocity gradients, leading to an underestimation of their rotational support. Other key observational limitations include spectral resolution and the signal-to-noise ratio (S/N) of the emission lines (e.g., Davies et al., 2011). At high S/N, velocity dispersions can be reliably measured down to one-third of the spectral resolution, but at low S/N, noise-driven line broadening inflates dispersion estimates, biasing $v/\sigma$ to lower values (e.g., Wisnioski et al., 2015; Übler et al., 2019).
@@
Given the impact of beam smearing and S/N limitations, the observed correlation between disk morphology and high $v/\sigma$ might be even stronger than what we found, with some dynamically cold disks being misclassified due to observational effects.
@@
We further investigated how effective radius and Sérsic index correlate with measured $v/\sigma$ ratios. To do this, we use the best-fit Sérsic parameters derived with statmorph on the sample that was matched in spatial resolution and rest-frame wavelength, as discussed in §5. The scatter plots in Figure 15 show $v/\sigma$ ratios as a function of $r_{\mathrm{eff}}$ and $n$. In terms of the Sérsic indices, we find no discernible trend since there is a clear clustering of both disks and non-disks below a threshold of $n=1.5$. In terms of the galaxy size, we detect a weak yet statistically significant positive correlation for disks between $v/\sigma$ and the effective radius, with a Spearman correlation coefficient ($\rho=0.26$, $p=0.004$). While the correlation is weak, it aligns with theoretical expectations and the broader connection between galaxy size and dynamical structure observed in both simulations and observational studies (e.g., Dekel et al., 2013; Fall & Romanowsky, 2013; Obreschkow & Glazebrook, 2014; Teklu et al., 2015; Posti et al., 2018; El-Badry et al., 2018; Tiley et al., 2021; Espejo Salcedo et al., 2025). Furthermore, larger disks are easier to visually identify in our classifications, making their structural properties more straightforward to analyze, as discussed in §3.2 and shown in Figure 7.
@@
Beyond the technical aspects of classification agreement, these results provide insight into the nature of galaxy evolution at these redshifts. The strong connection between disk morphology and high $v/\sigma$ suggests that a significant fraction of star-forming galaxies at cosmic noon are already dynamically settled and rotation-dominated. This is consistent with findings that a large fraction of galaxies at $z\sim 1-3$ exhibit regular rotation (e.g., Wisnioski et al., 2015; Johnson et al., 2018; Genzel et al., 2020; Tiley et al., 2021; Nestor Shachar et al., 2023). Moreover, we also find that a substantial fraction of these rotationally supported systems exhibit non-axisymmetric structures. Of the kinematically analyzed disks, 80 (64%) are classified as spirals, and 12 (10%) are classified as barred, with spiral structure, also identified in 10 of these. The median $v/\sigma$ values further support the idea that these structures form in dynamically cold disks, with spirals showing a median $v/\sigma\approx 7$ and bars a slightly lower median of $v/\sigma\approx 5$. The slightly lower rotational support in barred galaxies may reflect additional kinematic complexities introduced by bars, such as radial motions that enhance velocity dispersions.

### v2

@@ 4.4 Interacting and merger systems
The fraction of galaxies classified as interacting or merging systems shows a clear decline with cosmic time, decreasing from $0.14\pm 0.04$ in the higher redshift bin ($2.52.0<z<$ 2.5) to $0.06\pm 0.02$ in the lower-redshift bin ($1.30.8<z<$ 1.3). This trend aligns with the expected decrease in merger activity at lower redshifts, driven by declining galaxy densities and interaction rates (e.g., Conselice, 2007; Lin et al., 2008; Lotz et al., 2008; Rodriguez-Gomez et al., 2015). Despite our modest sample size, this decline remains consistent with theoretical and observational predictions in the framework of the hierarchical growth model in which major mergers become less frequent at later cosmic times.
@@ 5 Quantitative morphological metrics
We performed a quantitative analysis of the morphological properties of the sample to complement the visual classification. To achieve this, we employed several established nonparametric morphological indicators commonly utilized in morphological studies to characterize the structural features of the galaxies in our sample. These metrics include the concentration, asymmetry, and smoothness (CAS) statistics (Conselice, 2003), the multinode, intensity, and deviation (MID) statistics (Freeman et al., 2013), and the Gini and $M_{20}$ coefficients (Lotz et al., 2004). All measurements are quantified using statmorph (Rodriguez-Gomez et al., 2019). Here is a brief description of them.
@@
Figure 15: Notable parameter spaces where the combination of morphological metrics separate well different morphologies. In each panel, the red line corresponds to the line of best separation, which corresponds to the minimal contamination $\mathcal{C}_{\mathrm{min}}$ (defined in Appendix C), and the light red area indicates other possible line positions which achieve contamination within 0.01 of the minimal contamination, i.e., $C\in(\mathcal{C}_{\mathrm{min}},\mathcal{C}_{\mathrm{min}}+0.01)$. Left: Gini vs $M_{20}$ plane for galaxies with and without bulges. The line $G=-0.52M_{20}-0.4$, separates the main cluster of galaxies with bulges (71%) and those without bulges (74%). Middle: Gini vs $M_{20}$ plane for galaxies identified as interacting/mergers. The line $G=-0.29M_{20}+0.07$ (red) shows a systematic separation of the cluster of data points in this parameter space, with 50% of visually classified mergers and 93% of nonmergers correctly separated. For comparison, we show with the black dashed line the boundary line established by Lotz et al. (2008) and discuss the difference in the main text. Right: Gini vs asymmetry A plane for clumpy and nonclumpy galaxies. The separation line is given by $G=0.229A+0.448$, where 81.8% of clumpy galaxies are below the line and 64.1% of the nonclumpy galaxies are above.
MID: The multinode, intensity, and deviation (MID) statistics allow the identification of disturbed morphologies, such as mergers or interacting systems. Multinode (M) measures the presence of multiple peaks in the light distribution, indicating complex or multi-component structures. Intensity (I) assesses the brightness contrast between these peaks and their surroundings, highlighting areas of intense star formation or disruptions. Deviation (D) quantifies the departure of the light distribution from a smooth, symmetric model, further aiding in the detection of structural disturbances.
@@
To assess the robustness of the morphological indicators, we conducted a systematic visual inspection of diagnostic images produced by starmorph (e.g., Figure 4 in Rodriguez-Gomez et al., 2019). The reliability of each measurement was assessed based on the following criteria: How well the model replicates the galaxy’s light distribution (e.g., the agreement between the observed and model images); quality of the residual image, where well-fit galaxies exhibit minimally structured residuals and poor fits show significant leftover flux or systematic artefacts; convergence of fitted parameters, ensuring they fall within physically reasonable ranges; and absence of contamination from neighbouring sources or image artefacts that could bias measurements. The evaluation of these diagnostic images, along with assessments of segmentation maps and PSF matching, were used to ensure the accuracy of the nonparametric morphological measurements. After applying this filtering process, we identified a high-quality subset of 1,222 galaxies that met our reliability criteria. We adopted this subset for the results and analysis presented in this work.
@@ 5.1 Dependence on redshift and stellar mass
Figure 16: Top panel: Distribution of the asymmetry statistics (A in CAS) in our sample. The statistics show a clear systematic separation around value A$=0.23$. Maintaining a merger fraction of $\approx 10\%$, we obtained the vertical red line A$=0.332$ that separates 44% of mergers and 93% of nonmergers, with only a minor offset from the boundary line of A$=0.35$ introduced by Conselice (2003). Bottom three panels: Multinode, intensity, and deviation (MID) statistics. They reveal systematic differences in distributions between visually classified mergers and nonmergers, as expected, given their design to capture morphological features of disturbed systems.
Our findings also agree with trends reported in the literature, particularly in Yao et al. (2023) and Ren et al. (2024), which similarly noted stable morphological indicators across redshifts but a clear dependence on galaxy stellar mass. Figure 14 illustrates these comparisons, showing the mean values reported by Yao et al. (2023) for a F [^4] W-matched sample in the redshift range $30.8<z<$ 3, as well as the median values from Ren et al. (2024) for $30.5<z<$ 3. To enable a consistent comparison, we reverse-applied the observational corrections used in Ren et al. (2024) to their data, which accounted for effects such as noise and finite PSF resolution, aligning their values with our F [^4] W-matched dataset. In both cases, we found a strong agreement between the trends in our analysis and those reported by these studies.
@@ 5.2 Connection with visual classification
To use the extensive set of morphological parameters computed in this analysis, we systematically explored how different galaxy classifications separate across various parameter spaces derived from statmorph . Specifically, we identified parameter combinations that effectively distinguish galaxies with and without bulges, mergers and nonmergers, and clumpy versus nonclumpy galaxies. To quantitatively assess the separation between classifications, we computed a “separation efficiency” metric based on misclassification rates. The methodology for defining and calculating this metric, including the optimization of classification boundaries, is detailed in Appendix C, where we provide a full set of separation efficiencies for each classification. Given the wealth of possible parameter combinations, we highlight only the most informative and physically meaningful ones, prioritizing those that offer the clearest distinctions between morphological classes that have also been studied in the literature. However, all possible combinations, including those not well studied in other studies, are presented in Tables 4-6.
We emphasize that widely used morphological metrics, originally designed and quantified from HST imaging, must be carefully reassessed using JWST’s improved resolution, wavelength coverage, and depth. However, in this study, we focus on identifying broad, top-level trends in our sample rather than proposing precise recalibrations of these classification boundaries. A robust recalibration, which is beyond the scope of this work, should be carried out using larger statistical samples that fully leverage the growing wealth of high-quality JWST data, as well as the extensive galaxy catalogs from ongoing large surveys produced by Euclid, which will provide millions of galaxies at local and intermediate redshifts (e.g., Euclid Collaboration et al. 2025a, b).
Bulges: For galaxies visually identified as containing bulges, the Gini–$M_{20}$ plane provides one of the strongest separations between bulge-dominated and nonbulge galaxies (left panel of Figure 15). The optimal separation line, given by $G=-0.52M_{20}-0.41$, correctly separates 74% of galaxies without bulges, as shown by the clear systematic offset between the two populations in the figure. This offset aligns with previous studies (e.g., Conselice, 2003; Lotz et al., 2008), which demonstrated that nonparametric morphology indicators are effective in distinguishing bulge-dominated galaxies. The correlation between bulge presence, high Gini values, and low $M_{20}$ values supports the interpretation that bulge-dominated galaxies have centrally concentrated light distributions and less fragmented substructures, making them easier to identify in this parameter space. Specifically, low $M_{20}$ values indicate that a galaxy’s brightest pixels are concentrated near the center, which is characteristic of systems that contain prominent bulges.
@@
Mergers: Several parameter combinations effectively distinguish interacting and merging galaxies, including the Gini–$M_{20}$ plane, MID statistics, and CAS parameters. The full separation efficiencies for mergers and nonmergers are presented in Table 5 in the Appendix.
@@
In the Gini–$M_{20}$ plane, we define a separation line that maintains a merger fraction of $\approx 10\%$, matching both our measured fraction and the findings of Lotz et al. (2008). The resulting threshold, given by $G=-0.29M_{20}+0.07$, correctly captures 93% of nonmergers (middle panel of Figure 15). For comparison, we show the boundary line estimated by a dedicated study of mergers and morphology Lotz et al. (2008), which analyzed $\approx 3,000$ galaxies with HST imaging at $1.20.2<z<$ 1.2 and found a separation boundary of $G=-0.14M_{20}+0.33$. We found notable differences between our classification and theirs, particularly a 0.15 slope difference in the separation line. Moreover, our separation achieves a contamination value of 0.57, while adopting the Lotz et al. boundary in our dataset would result in a higher contamination value of 0.74. The differences in our boundary lines likely stem from differences in sample selection, imaging depth, rest-frame wavelength, spatial resolution, and classification objectives. Whereas Lotz et al. optimized their boundary specifically for merger detection, we focused on a wider range of morphologies. Despite these differences, the overall separation trends remain similar, reinforcing that Gini vs $M_{20}$ remains a useful discriminator and is effective, at least in a global sense, within our sample.
Figure 17: Scatter plots of $v_{\mathrm{rot}}(r_{\mathrm{eff}})/\sigma_{0}$ as a function of effective radius (left) and Sérsic index (right) for disk (orange circles) and nondisk (blue squares) galaxies. The histograms above each panel show the normalized distributions of effective radius and Sérsic index, while the histogram on the right represents the distribution of $v_{\mathrm{rot}}(r_{\mathrm{eff}})/\sigma_{0}$. The red horizontal lines indicate reference values at $v_{\mathrm{rot}}(r_{\mathrm{eff}})/\sigma_{0}=1,3,$ and $10$.
@@
Among the CAS statistics, the asymmetry parameter provides another useful merger discriminator. Figure 16 (top panel) illustrates the separation, where visually identified mergers are skewed toward higher asymmetry values, while nonmergers peak at lower asymmetries. The separation threshold that maintains a merger fraction of 10% closely aligns with the empirical merger criterion proposed by Conselice (2003) (dashed black), reinforcing the utility of asymmetry as a merger diagnostic.
@@
Similarly, the MID statistics also effectively separate mergers from nonmergers. As shown in the bottom panels of Figure 16, the multinode (M), intensity (I), and deviation (D) distributions exhibit systematic differences between the two populations. Interacting systems consistently show higher values across all three statistics, reinforcing the idea that MID metrics capture morphological disturbances and asymmetries associated with mergers and their irregular light distributions.
@@
Clumps: The Gini coefficient is particularly effective in distinguishing clumpy from nonclumpy galaxies, especially when combined with other morphological metrics. Among these, the Gini–A (asymmetry) plane provides the clearest separation (right panel of Figure 15), highlighting the strong correlation between light concentration and structural irregularity in clumpy systems. The complete set of separation efficiencies for clumpy galaxies is provided in Table 6 in the Appendix.
@@ 6 Morpho-kinematic Correlations
For this discussion, we focus on global morpho-kinematic trends on a population-wide angle rather than detailed substructures, examining how the identified morphologies relate to dominant dynamical support as measured by $v/\sigma$. We used the homogenized $v_{\mathrm{rot}}(r_{\mathrm{eff}})$ and $\sigma_{0}$ values obtained from various subsamples outlined in Sect. 2.3. Our analysis reveals a notable distinction in the dynamical properties of galaxies classified as disks versus nondisks. For the 155 galaxies that have kinematics measurements, 125 (fraction $\approx 0.81$) were visually classified as disks, while 30 ($\approx 0.19$) were not identified as disks.
@@
The histogram on the right of Figure 17 presents the $v/\sigma$ distributions for visually classified disks and nondisks. A high $v/\sigma$ ratio suggests a dynamically cold, rotation-dominated system typical of disks, while lower values indicate dynamically hot, dispersion-dominated systems more consistent with spheroidal geometries or disturbed morphologies. We include horizontal lines at $v/\sigma=1$, 3, and 10—commonly adopted thresholds in the literature corresponding to varying degrees of rotational support (e.g., Genzel et al., 2006; Kassin et al., 2012; Wisnioski et al., 2015; Wisnioski et al., 2019; Stott et al., 2016; Tiley et al., 2019; Gillman et al., 2020; Förster Schreiber & Wuyts, 2020).
@@
To quantify the difference between the two populations, we performed a Kolmogorov–Smirnov (KS) test, which yields a statistic of 0.39 and a $p$-value of $9.4\times 10^{-4}$, confirming that the $v/\sigma$ distributions of disks and nondisks are significantly different. Specifically, among the 125 visually classified disk galaxies, 88 (fraction $\approx 0.70$) exhibit $v/\sigma>3$, indicating strong rotational support. The remaining 37 ($\approx 0.30$) fall below this threshold, suggesting either weaker rotational support or the presence of noncircular motions (pure turbulence but also radial motions contributing to higher $\sigma_{0}$ in modeling of the typically seeing-limited IFU data included here). For nondisk galaxies, the trend is reversed: only 12 galaxies ($\approx 0.40$) have $v/\sigma>3$, while the majority, 18 galaxies ($\approx 0.60$), have $v/\sigma<3$. This distribution reflects the expectation that nondisk galaxies exhibit lower rotational support, as their dynamics are often dominated by random motions rather than ordered rotation. In the case of marginal rotational support of $v/\sigma=1$, only three galaxies identified as disks and five from the other categories are below this boundary.
@@
The apparent correlation between disk morphology and high $v/\sigma$ is reassuring, yet it is important to consider observational limitations in the kinematic measurements that could affect such correlation. One such effect is beam smearing, which artificially enhances velocity dispersions in smaller galaxies (e.g., Newman et al., 2013). In particular, compact galaxies at high redshift may exhibit inflated velocity dispersions due to unresolved velocity gradients, leading to an underestimation of their rotational support. Other key observational limitations include spectral resolution and the signal-to-noise ratio (S/N) of the emission lines (e.g., Davies et al., 2011). At high S/N, velocity dispersions can be reliably measured down to one-third of the spectral resolution, but at low S/N, noise-driven line broadening inflates dispersion estimates, biasing $v/\sigma$ to lower values (e.g., Wisnioski et al., 2015; Übler et al., 2019).
@@
Given the impact of beam smearing and S/N limitations, the observed correlation between disk morphology and high $v/\sigma$ might be even stronger than what we found, since some dynamically cold disks could be misclassified due to observational effects.
@@
We further investigated how effective radius and Sérsic index correlate with measured $v/\sigma$ ratios. To do this, we used the best-fit Sérsic parameters derived with statmorph on the sample that was matched in spatial resolution and rest-frame wavelength, as discussed in Sect. 5. The scatter plots in Figure 17 show $v/\sigma$ ratios as a function of $r_{\mathrm{eff}}$ and $n$. In terms of the Sérsic indices, we found no discernible trend since there is a clear clustering of both disks and nondisks below a threshold of $n=1.5$. In terms of the galaxy size, we found a weak yet statistically significant positive correlation for disks between $v/\sigma$ and the effective radius, with a Spearman correlation coefficient ($\rho=0.26$, $p=0.004$). While the correlation is weak, it aligns with theoretical expectations and the broader connection between galaxy size and dynamical structure observed in both simulations and observational studies (e.g., Dekel et al., 2013; Fall & Romanowsky, 2013; Obreschkow & Glazebrook, 2014; Teklu et al., 2015; Posti et al., 2018; El-Badry et al., 2018; Tiley et al., 2021; Espejo Salcedo et al., 2022, 2025). Furthermore, larger disks are easier to visually identify in our classifications, making their structural properties more straightforward to analyze, as discussed in Sect. 3.2 and shown in Figure 8.
@@
Beyond the technical aspects of classification agreement, these results provide insight into the nature of galaxy evolution at these redshifts. The observed connection between disk morphology and high $v/\sigma$ suggests that a significant fraction of star-forming galaxies at cosmic noon are already dynamically settled and rotation-dominated. This is consistent with findings that a large fraction of galaxies at $z\sim 1-3$ exhibit regular rotation (e.g., Wisnioski et al., 2015; Johnson et al., 2018; Genzel et al., 2020; Tiley et al., 2021; Nestor Shachar et al., 2023). Moreover, we also found that a substantial fraction of these rotationally supported systems exhibit non-axisymmetric structures. Of the kinematically analyzed disks, 80 (fraction 0.64) are classified as spirals, and 12 (0.10) are classified as barred, ten of which also have spiral structure. The median $v/\sigma$ values further support the idea that these structures form in dynamically cold disks: spirals show a median $v/\sigma\approx 7$ and bars show a slightly lower value of $v/\sigma\approx 5$. This modest drop in rotational support for barred galaxies may reflect added kinematic complexity associated with bar-driven secular evolution, in line with recent discussions on angular momentum redistribution and dynamical stability in barred disk galaxies (e.g., Romeo et al., 2023).